OpenStudy (anonymous):

17) During a store sale, the regular price for jeans is discounted by $7. There is a state sales tax of 3%, and the $7 discount is applied before the sales tax is calculated. a. Write an expression that shows the regular price r of a jeans minus the $7 discount. b. Write a rule for the function p(r) that expresses the final price p of jeans with the discount applied and sales tax added. c. How much would you pay during the sale for jeans regularly priced at $24.99?

OpenStudy (dinmosiaren):

Hey ellie

OpenStudy (dinmosiaren):

no im not...and no i cant help with this

OpenStudy (anonymous):

(a) r-7 (b)net price p=(r-7)+(r-7)*3% (c) price after discount=24.99-7=17.99 sale tax=17.99*3%=.5397=.54 nearly you pay=17.99+.54=$18.53

OpenStudy (anonymous):

Thank you so much!